OpenStudy (anonymous):

Use continuity to evaluate the limit, as the limit of x approaches pi of 3sin(x+sinx)

OpenStudy (science0229):

\[\lim_{x \rightarrow \pi}3\sin (x+\sin x)\]

OpenStudy (hba):

Use L'Hopital

OpenStudy (anonymous):

@hba , l'hopital? what's that

OpenStudy (hba):

Well you just diffrentiate it so that you don't end in a 0/0 form.This one looks simple so you'll have to just apply limit here.

OpenStudy (dumbcow):

hmm no just evaluate limit at pi

OpenStudy (science0229):

The function is continuous on all values of x.

OpenStudy (science0229):

That means that the limit of every x value is same as the y-value. In other words, \[\lim_{x \rightarrow a}f(x)=f(a)\] where a is a x-value on f(x)

OpenStudy (science0229):

So, \[\lim_{x \rightarrow \pi}3\sin(x+\sin x)=f(\pi)\]

OpenStudy (science0229):

Evaluate f(pi) and you'll get 0.